Creativeone15:

Jason paid $15.50 for 3 slices of pizza and 2 burgers. Susan paid $20 for 1 slice of pizza and 4 burgers. Fill in the following:(Do NOT include a $ in your answer) Each slice of pizza costs $ ___________________________ . Each burger costs $ _________________________

Vocaloid:

Moved to Mathematics anyway, you want to create a system and solve it let p = # of pizzas and b = # of burgers Jason's order: 3p + 2b = 15.50 Susan's order: p + 4b = 20 solve for p and b using elimination or substitution

Creativeone15:

ok

Vocaloid:

still there? have you made any progress so far?

Vocaloid:

I'm going to close this for now but if you ever want to get back to this just tag me @Vocaloid

Creativeone15:

ok i have more..........

Creativeone15:

@Vocaloid

Creativeone15:

Select the correct choices that belong in the blank for the system of equations shown below. 24x - 39 = 165 -6x +13y = -51 In order to solve this system by elimination, Stacy ____________ by ______. Then, when she adds the equations together, the x terms will cancel. Question 4 options: multiplies the second equation; -4 multiplies the first equation; 4 multiplies the second equation; 4 multiplies the first equation; -4

Vocaloid:

the second equation has a -6x and the first one has a 24x term what do we have to multiply -6x by to turn it into -24x? (so that the terms can cancel out through addition)